How much water must be evaporated from 32 ounces of a 5% salt solution to make a 8% salt solution?
Answer by ptaylor(2198)

Let x=amount of water that needs to be evaporated
Now we know that the amount of pure salt that we have before we evaporate the water has to equal the amount of pure salt that we have after we evaporate the water.
Amount of pure salt before we evaporate=0.05*32
Amount of pure salt after we evaporate= 0.08(32-x)
So, our equation to solve is:
0.05*32=0.08(32-x) simplify
1.6=2.56-0.08x subtract 2.56 from each side
-0.08x=-0.96
x=12 oz-----amount of water that needs to be evaporated
CK
32*0.05=20*0.08
1.6=1.6
Hope this helps---ptaylor